OpenStudy (anonymous):

Can someone check my work and see where I went wrong ? Find an equation of line tangent to the graph of y=1-8x-9x^2 at the point (2,-51). Answer:y=________ Point-slope formula y-(-51)=m(x-2) f'(x)= -18x-8 f'(2)=-18(2)-8=-44 y-(-51)=-44(x-2) y=-44x+139 and I don't know where I went wrong

OpenStudy (psymon):

You added 51 instead of subtracted y -(-51) = -44(x-2) y + 51 = -44x + 88 y = -44x + 37

OpenStudy (anonymous):

OMG really, I didn't catch that thanks ! :D

OpenStudy (psymon):

Yeah, np, lol.